OpenStudy (anonymous):

In the elimination method . When multiplying and equation to get two opposites, do you multiply after the equal sign? for example. 4(5x-4y=-19 would it be 20x-16y=-19 still or -76

OpenStudy (anonymous):

The Latter. You also multiply -19 by 4

OpenStudy (phi):

if you have an equation, example 1+1 = 2 it is still an equation (both sides are equal to each other) if you double both sides: 2*(1+1) = 2*2 or if you multiplied both sides by 4 or by x... you have to multiply both sides of the = to keep the equation true